OpenStudy (anonymous):

What is the amplitude, period, and phase shift of f(x) = -4 sin(2x + π) - 5?

OpenStudy (anonymous):

amplitude = -4; period = 2π; phase shift: x = -pi/2 amplitude = -4; period = π; phase shift: x = pi/2 amplitude = 4; period = π; phase shift: x = -pi/2 amplitude = 4; period = 2π; phase shift: x = pi/2

OpenStudy (anonymous):

@ganeshie8 @Calliope @random231 @ranga

random231 (random231):

any ideas?

OpenStudy (anonymous):

Either first or last i know its a shift to the right?

OpenStudy (anonymous):

@random231

random231 (random231):

well can amplitude be negative?

OpenStudy (anonymous):

No? Arent there certain limitations on Amplitude on graphs, and being negative is one? @random231

random231 (random231):

well amplitude is the measure of the maximum lateral displacement of a wave. so in my opinion an amplitude cannot have an negative sign!

random231 (random231):

@Mashy ! wave equation help!

OpenStudy (anonymous):

I am so terrible at math! Lol

random231 (random231):

lol actually im not very sure! ive asked @Mashy to come here and im sure he can help.probably he is a bit busy!

random231 (random231):

but im pretty sure that amplitude wont be -ve

OpenStudy (anonymous):

okay. so that narrows it down to one or the other answer

OpenStudy (anonymous):

sin( pi + x) = -sin(x) so u can reduce that equation to f(x) = 4sin ( 2x) + 5

OpenStudy (anonymous):

sorry i mean - 5*

OpenStudy (anonymous):

Okay. So then its C?

OpenStudy (anonymous):

well.. i think so.. i don't get the phase shift part though.. :D

OpenStudy (anonymous):

Okay let me see if its right on the quiz...

OpenStudy (anonymous):

Yes, its C! Thanks guys! @Mashy @random231

OpenStudy (anonymous):

np!

random231 (random231):

np!